OpenStudy (anonymous):

Person 1 can do a certain job in twenty-one minutes, and person 2 can do the same job in twenty-eight minutes. When completing the job together, which expression would be used to represent the amount of work done by person 1? a. 21 b. 1/21 c. x/21 d. 21x

OpenStudy (anonymous):

b? mnot sure..

OpenStudy (shane_b):

To set up the problem you add up the rates like this:\[\frac{1}{21}+\frac{1}{28}=\frac{1}{total\space if\space done\space together}\]

OpenStudy (shane_b):

So the left fraction will be would be the rate of person 1.

OpenStudy (shane_b):

IF the question provided the rates of each person and you wanted to know how fast they could do the work together you could also use this shortcut:\[\frac{r_1r_2}{r_1+r_2}=total\space if\space done\space together\]

OpenStudy (anonymous):

so b is wrong?

OpenStudy (shane_b):

No...b is right.